Determine whether the vector (9,0,9) is a linear combination of the vectors


(1,1,0), (1,0,1), (2,1,1) and (0,1, −1).

1
Expert's answer
2022-04-08T03:58:38-0400

There is an obvious linear combination:

"(9,0,9)=0\cdot(1,1,0)+9\cdot(1,0,1)+0\cdot(2,1,1)+0\cdot(0,1,-1)."